OpenStudy (anonymous):

right triangle abc is graphed on the coordinate plane and has vertices at a(-1,3), b(0,5), and c(4,3). what is the measure of

OpenStudy (anonymous):

OKay, so I know this: \[\sin (C)=\frac{ \sqrt{5} }{5}\] But i dont know how to get the angle.